20) Six tubes are in a rack.. You mix 1ml of cell culture with 0.5 ml of water in tube #1, transfer 0.1 ml of tube #1 to tube #2 containing 0.9ml of water and mix well and the same procedure (0.1 ml of culture +0.9 ml of water) is repeated for the remaining 4 tubes. After plating 0.1 ml of the last dilution tube #6, you obtain 100 colonies. what is the cell density (cfu/ml) of the original cell culture?
Answer #1

Answer: 3 X 107 cfu/mL

Explanation:

Initial volume = 1 mL
Initial concentration = X
Tube 1 = 1 ml + 0.5 mL
Tube 1 concentration = X/3

Tube 2 = 0.1 ml + 0.9 mL
Tube 2 concentration = X/3 X 1/10

Tube 3 = 0.1 ml + 0.9 mL
Tube 3 concentration = X/3 X 1/10 X 1/10

Tube 4 = 0.1 ml + 0.9 mL
Tube 4 concentration = X/3 X 1/10 X 1/10 X 1/10

Tube 5 = 0.1 ml + 0.9 mL
Tube 5 concentration = X/3 X 1/10 X 1/10 X 1/10 X 1/10

Tube 6 = 0.1 ml + 0.9 mL
Tube 6 concentration = X/3 X 1/10 X 1/10 X 1/10 X 1/10 X 1/10
Tube 6 concentration = X/3 X 1/100000
Tube 6 concentration = X/300000
X/300000 = 100 colonies

Original sample concentration, X = 100*300000 = 30000000 cfu/mL
= 3 X 107 cfu/mL